The Downtown Abbey Brewery is considering a new four-year expansion project that requires an initial fixed asset investment of $210,000. The fixed asset will be depreciated straight-line to zero over its four-year life, after which time it will be worthless. The project is estimated to generate $48,000 in annual sales, with costs of $31,000. If the tax rate is 34 percent, what is the OCF for this project?
Answer #1

Annual depreciation = 210,000 / 4 = 52,500

OCF = (Sales - costs - depreciation)(1 - tax) + depreciation

OCF = (48,000 - 31,000 - 52,500)(1 - 0.34) + 52,500

OCF = $29,070